solves the following linear equation:
5x over 3 + 2 = 7x + 1 over 4

Answer by unlockmath(1688)

Hello,
Let's rewrite this 5x over 3 + 2 = 7x + 1 over 4 as:
5x/3 + 2 = 7x + 1/4 Now let's get rid of the fractions by multiplying by 12 to get:
20x+24=84x+3 Now subtract 3 and 20x from both sides to give us:
21=64x Divide both sides by 64 to get:
x=21/64
Check it out by plugging in the answer.
RJ
Check out a new book at
www.math-unlock.com